# Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning

### v1

@@ 6.2 Impact of Prompting Strategies
Limitations of Chain-of-Thought. CoT yields mixed results, slightly improving borderline-case identification. For instance, GPT-4o on DBDC5 English F1(B) improves from 85.9% to 86.5%. However, this occasionally degrades performance on structured datasets like BETOLD, indicating an optimal reasoning complexity is dataset-dependent. This shows the importance of calibrating the number of reasoning steps: dialogues with short turns and domain-specific structure can be more efficiently parsed with concise prompts than with extensive “think step-by-step” sequences.
@@
Challenges with AR and CL+AR on BETOLD. While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD. Results for BETOLD (highlighted in grey in Table 2) consistently showed lower performance. We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances. This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue. Through comprehensive error analysis, we observed that models frequently did not fully adhere to provided instructions. Common issues included either failing to generate any analogous examples—opting instead to solve the original dialogue directly—or producing overly brief summaries of analogous conversations that did not align well with the dialogue under evaluation. As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions. This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns). Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3. Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
6.3 Calibration and Confidence Analysis
@@
Superior calibration (low MSE) indicates alignment between model confidence and human correctness. Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard). Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard). Meanwhile, our fine-tuned 8B monitor yields an MSE of 4.9, which is slightly higher but still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
6.4 Practical Costs and Sustainability
@@
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[46]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 317 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 4,067 times higher [47]. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
@@
In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency—each dialogue inference completes in under 0.5 seconds. This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance. Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to slash costs by more than half compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations). Such selective invocation lowers operational costs and aligns with sustainability objectives. Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S. households [48]. By limiting resource-intensive model invocations to cases of genuine ambiguity or critical dialogue breakdowns, our approach effectively balances performance, cost-efficiency, and environmental sustainability.
7 Conclusion
@@
Research and development in conversational AI robustness is rapidly advancing on multiple fronts. Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively. By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation. While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains. Short, curated exemplars typically offer a good trade-off between performance and token usage. The cost and energy savings are considerable, especially in frequent-turn applications. Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness. Our experiments reveal that carefully structured prompts—where the model justifies and then decides—yield lower calibration error. For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability. While closed-source models still lead in absolute accuracy, the performance gap is closing—especially once small and midsize open-source models are carefully tuned. Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
8 Limitations
@@ Appendix A Cost-Savings Example on AWS Bedrock
To illustrate how real-time monitoring reduces expenses, we assume a 15-turn conversation where a user interacts primarily with Llama-3.3 70B but escalates to Llama-3.1 405B on 10% of turns if a breakdown is detected. Table 3 shows prices [^3] https://aws.amazon.com/bedrock/pricing/ per 1,000 tokens for Meta’s Llama-3.1 family.
Table 3: Selected AWS Bedrock prices for Llama 3.1.
Model Price / 1K Input Price / 1K Output
@@
Llama-3.1 405B \$0.0024 \$0.0024
Assumptions
  • On average, each turn has 160 characters444www.twilio.com/docs/glossary/what-sms-character-limit (approximately 40 tokens) in the user prompt plus 40 tokens in the assistant’s reply, for 80 tokens total per turn.
  • At turn $i$, the LLM sees the entire conversation so far—so the token count accumulates.
  • Only 10% of turns require escalation to the 405B model.
Calculating Total Tokens Over 15 Turns
Since each turn $i$ includes all previous turns in the prompt, the approximate token usage over 15 turns:
$80\times\sum_{i=1}^{15}i=80\times\frac{15\times(15+1)}{2}=80\times 120=9600{}\text{tokens}$
Cost for Always Using 405B
  • Price: \$0.0048 per 1,000 tokens.
  • Total: $9.6\times\$0.0048\approx\$0.04608$
Cost for Always Using 70B
  • Price: \$0.00144 per 1,000 tokens.
  • Total: $9.6\times\$0.00144\approx\$0.01382$
Selective Escalation (70B + 10% 405B)
Assuming 10% of the total tokens trigger a breakdown and invoke the 405B model, while the remaining 90% stay on 70B:
  • 70B portion (8640 tokens): $8.64\,\text{k}\times\$0.00144=\$0.01244$
  • 405B portion (960 tokens): $0.96\,\text{k}\times\$0.0048=\$0.00460$
  • Combined: \$$\$0.01244+\$0.00460\approx\$0.01704$
Adding a Monitor for Real-time Checks
Every turn is first checked by a monitor:
  1. 1.
    Monitoring Cost:
    • 9,600 tokens at \$0.00044 per 1,000 tokens
    • $9.6\,\text{k}\times\$0.00044\approx\$0.00422$
  2. 2.
    Main Dialogue:
    • Still 90% on 70B and 10% on 405B \$$\approx\$0.01704$
  3. 3.
    Total: \$$\$0.00422+\$0.01704=\$0.02126$
Hence, even with monitoring overhead, escalating only 10% of turns to the 405B model cuts costs by more than half compared to always using Llama-3.1 405B.

### v2

@@ VI-B Impact of Prompting Strategies
Limitations of Chain-of-Thought. CoT yields mixed results, slightly improving borderline-case identification. For instance, GPT-4o on DBDC5 English F1(B) improves from 85.9% to 86.5%. However, this occasionally degrades performance on structured datasets like BETOLD, indicating that an optimal reasoning complexity is dataset-dependent. This highlights the importance of tailoring reasoning complexity to the task: dialogues with short turns or highly structured, domain-specific content may be more effectively processed using concise prompts rather than elaborate ‘think step-by-step’ sequences.
@@
Challenges with AR and CL+AR on BETOLD. While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD. Results for BETOLD (highlighted in grey in Table II) consistently showed lower performance. We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances. This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue. Through comprehensive error analysis, we observed that models frequently did not fully adhere to the provided instructions. Common issues included either failing to generate any analogous examples, opting instead to solve the original dialogue directly, or producing overly brief analogous summaries that poorly aligned with the dialogue under evaluation. As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions. This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns). Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3. Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
VI-C Calibration and Confidence Analysis
@@
Superior calibration (low MSE) indicates alignment between model confidence and human correctness. Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard). Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard). Meanwhile, our fine-tuned 8B Dialogue Disruption Monitor yields an MSE of 4.9. While slightly higher, this result still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
VI-D Practical Costs and Sustainability
@@
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[45]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [46]. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
@@
In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency (each dialogue inference completes in under half a second). This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance. Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to reduce costs by approximately 54% compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations). Such selective invocation lowers operational costs and aligns with sustainability objectives. Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S. households [47]. By limiting resource-intensive model invocations to instances of suspected critical dialogue breakdowns or high uncertainty, our approach effectively balances performance, cost-efficiency, and sustainability.
VII Conclusion
@@
Research and development in conversational AI robustness is rapidly advancing on multiple fronts. Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively. By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation. While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains. Short, curated exemplars typically offer a good trade-off between performance and token usage. The cost and energy savings are considerable, especially in frequent-turn applications. Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness. Our experiments reveal that carefully structured prompts, where the model justifies and then decides, yield lower calibration error. For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability. While closed-source models still lead in absolute accuracy, the performance gap is closing, especially once small open-source models are carefully tuned. Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
VIII Limitations
@@ Appendix A Cost-Savings Example on AWS Bedrock
This section quantifies the benefit of selective model escalation. We consider a 15-turn dialogue that normally runs on Llama-3.1 70B. When a real-time monitor detects a potential breakdown, the request is re-issued to the larger Llama-3.1 405B. Empirically, such escalations are required on roughly $10\%$ of turns.
Table III lists on-demand prices [^3] https://aws.amazon.com/bedrock/pricing/ for Meta’s Llama-3.1 family. Bedrock bills input and output tokens separately, so the effective cost per 1k tokens is the sum of the two columns.
TABLE III: Representative AWS Bedrock prices for Llama-3.1 (May 2025).
Model Input (per 1k) Output (per 1k)
@@
Llama-3.1 405B \$0.00240 \$0.00240
A-A Token Budget for 15 Turns
Each turn contributes $\approx\!40$ tokens in the user prompt and another $40$ tokens in the reply, for $80$ new tokens per turn. Because the entire history is sent at every step, turn $i$ carries $80\,i$ tokens. Over 15 turns:
Total tokens= $\text{Total tokens}\;=\;80\sum_{i=1}^{15}i\;=\;80\times\frac{15\times 16}{2}\;=\;9\,600\text{ tokens.}$.
A-B Baseline Costs
  • Always 405B  $\;9.6\times\$0.00480=\$0.0461$
  • Always 70B  $\;9.6\times\$0.00144=\$0.0138$
A-C Selective Escalation (70B + 10% 405B)
[…]
Total \$$\approx\$0.01705.$
A-D Adding a Lightweight Monitor
The monitor itself runs on Llama-3.1 8B and processes every turn:
[…]
Putting it all together:
\$$\$0.00422\;(\text{monitor})\;+\;\$0.01705\;(\text{dialogue})\;=\;\$0.02127.$
Selective escalation with monitoring cuts cost by $1-\tfrac{0.02127}{0.0461}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.